Barb is going to cover a rectangular area 8 feet by 10 feet with rectangular paving blocks that are 4 inches by 8 inches by 2 inches to make a flat patio. What is the minimum number of paving blocks she will need if all the paving blocks will face the same direction? Note: Barb will not cut any of the paving blocks.
Answer by tomlaidlaw(14)

Area of paving block = 8 x 4 = 32 in.^2
patio calcs:
8 ft.*12=96 inches
10 ft.*12=120 inches
Area of patio = 120*96 = 11520 in.^2
Number of pavers = 11520/32 = 360
Tom Laidlaw